OpenStudy (anonymous):

Solve: 91x  = 189 mod 231

OpenStudy (anonymous):

first find the inverse of 91 mod 231 using the euclidean algorithm

OpenStudy (anonymous):

scratch that, 91 and 231 are not relatively prime, they are both divisible by 7

OpenStudy (anonymous):

so divide everything by 7 and start with \[13x\equiv 27(\text { mod }33)\]

OpenStudy (anonymous):

we can find the inverse of 13 using the euclidean algorithm

OpenStudy (anonymous):

\[33-2\times 13+7\] \[13=7+6\] \[7=6+1\] work backwards to solve for 1

OpenStudy (anonymous):

first line should be \[33=2\times13+7\]

OpenStudy (anonymous):

work backwards and get \[1=7-6\] \[1=7-(13-7)=2\times 7-13\] \[1=2(33-2\times 13)-13=2\times 33-5\times 13\]

OpenStudy (anonymous):

so the "bezou coefficients" are 2 and -5, therefore -5 is the inverse of 13 mod 33 gotta run, hope you can finish from there answer is 30

OpenStudy (anonymous):

cook a bunch for me!!!